i am not sure i understand how to do this
A = 1/2bh
A = 1/2 (2.75)(8.25)

Answer by Nate(3500)

A = (1/2)(2.75)(8.25) = (1/2)(11/4)(33/4) = 363/32 = 11 11/32 = 11.34375